OpenStudy (anonymous):

3+logbase2(2x-5)=log(5x-7)

OpenStudy (king):

pleae post with equation editor

OpenStudy (king):

fr log 5x-7 what is the base?is base given?

OpenStudy (lgbasallote):

this log(5x-7)...are you sure it doesn't have a base 2? Because this problem seems hard the way it appears...

OpenStudy (anonymous):

\[\log_28+\log_2(2x-5) = \log_2(5x-7)\]

OpenStudy (king):

is the answer x=3?

OpenStudy (anonymous):

the same questions as lgbsallote

OpenStudy (king):

nutty is it x=3?

OpenStudy (anonymous):

for log (5x-7) the base is 2

OpenStudy (king):

then answer is x=3?u want solution?

OpenStudy (anonymous):

I dont' think so

OpenStudy (king):

substitute and see it is rite victorarana.....!!!!!!!!!!!!!!!

OpenStudy (lgbasallote):

then the solution is to put all the logs on one side. it becomes logbase 2 (2x-5)/(5x-7) = -3. 2^-3 = 2x - 5/5x - 7. 1/8 = (2x-5)/(5x-7). cross multiply. 5x - 7 = 16x - 40. 11x = 33. x = 3. King is right

OpenStudy (king):

thnx igbasallote!!!!!!

OpenStudy (anonymous):

let me think

OpenStudy (anonymous):

can i hve the solution plzzz

OpenStudy (anonymous):

yes it's not correct

OpenStudy (anonymous):

\[3 = \log_2\frac{5x-7}{2x-5}\]

OpenStudy (anonymous):

so \[2^3 = \frac{5x-7}{2x-5}\]

OpenStudy (king):

ya

OpenStudy (anonymous):

solving for x \[x=\frac{33}{14}\]

OpenStudy (anonymous):

are you agree? King?

OpenStudy (king):

no u hav solved wrong

OpenStudy (anonymous):

mm you're right it was 33/11 = 3

OpenStudy (king):

ya gud!

OpenStudy (anonymous):

sooo wats da answer knw cauze am soo confused

OpenStudy (king):

answer is x=3

OpenStudy (king):

ok?

OpenStudy (anonymous):

oky